# Scaling Verification Can Be More Effective than Scaling Policy Learning for Vision-Language-Action Alignment

### v1

@@ Abstract
The long-standing vision of general-purpose robots hinges on their ability to understand and act upon natural language instructions. Vision-Language-Action (VLA) models have made remarkable progress toward this goal, yet their generated actions can still misalign with the given instructions. In this paper, we investigate test-time verification as a means to shrink the “intention-action gap.” We first characterize the test-time scaling law for embodied instruction following and demonstrate that jointly scaling the number of rephrased instructions and generated actions greatly increases test-time sample diversity, often recovering correct actions more efficiently than scaling each dimension independently. To capitalize on these scaling laws, we present CoVer, a contrastive verifier for vision–language–action alignment, and show that our architecture scales gracefully with additional computational resources and data. We then introduce “boot-time compute” and a hierarchical verification inference pipeline for VLAs. At deployment, our framework precomputes a diverse set of rephrased instructions from a Vision-Language-Model (VLM), repeatedly generates action candidates for each instruction, and then uses a verifier to select the optimal high-level prompt and low-level action chunks. Compared to scaling policy pre-training on the same data, our verification approach yields 22% gains in-distribution and 13% out-of-distribution on the SIMPLER benchmark, with a further 45% improvement in real-world experiments. On the PolaRiS benchmark, CoVer achieves 14% gains in task progress and 9% in success rate.
@@ 1 Introduction
Figure 2: Test-Time Scaling Law for Embodied Instruction Following. Compared to prior methods that obtain diverse actions through repeated sampling [15] or Gaussian perturbations [12], we find that instruction rephrasing produces a broader set of action candidates, leading to improved recovery of the correct action. Furthermore, a hybrid test-time scaling strategy that increases both the number of rephrases and the number of sampled actions per rephrase is more effective than either strategy alone. We characterize each sampling approach using a power law, where the logarithm of oracle action error $e$ is a function of the number of action candidates $k$: $\log(e)\approx\log(a)+b\cdot\log(k)$.
@@ 4.1 Hierarchical Prompt-Action Optimization
Figure 3: Overview of CoVer Training Strategy. CoVer learns a joint embedding space aligning visual observations, language instructions, and robot actions through contrastive pre-training. Image and text encoders extract task-relevant visual–linguistic features, which are fused into text-aware visual representations. An action encoder maps action sequences into the same embedding space, enabling cross-modal alignment between instructions and executed behaviors.
@@ 5.4 Simulation Evaluation Results
Table 1: PolaRiS benchmark evaluation. Mean task progress and success rate ($\pm$ standard deviation) across 50 episodes and 3 seeds on three PolaRiS environments. $\pi_{0.5}$ serves as the stronger base VLA policy. $\pi_{0.5}$ + CoVer consistently improves performance across all tasks, achieving a 13.9% gain in task progress and a 9.3% increase in success rate on average, demonstrating that hierarchical verification complements stronger base models.
@@ 5.6 Latency Analysis and Optimizations
Figure 8: Real-World Evaluation Results. $\pi_{0}$ + CoVer outperforms the baseline policy $\pi_{0}$ (rephrase), achieving on average 45% performance gain in terms of task success rate compared to the baseline policy.
@@ 6 Conclusion
In this paper, we present CoVer, a novel contrastive-based verifier and hierarchical test-time scaling framework that bridges the “intention–action gap” for generalist robot policies. CoVer achieves substantial performance improvements across both simulated and real-world settings, particularly under out-of-distribution conditions. Our findings demonstrate that allocating compute to reasoning and verification at deployment can be more effective than scaling policy training alone, providing a promising direction for robust policy deployment in the real world. While our study focuses on applying the verifier for test-time scaling, the same design and principle can extend beyond inference optimizations such as post-training with reinforcement learning or run-time monitoring. Future work could also explore more efficient architectures for both the base policy and verifier to further reduce latency and enable broader use of test-time scaling in real-world robotic settings.
@@ 7 Acknowledgments
We thank the members of the Stanford Autonomous Systems Lab, Scaling Intelligence Lab, and IRIS Lab for their constructive feedback and informative discussions. We gratefully acknowledge the support of DARPA; NASA ULI; Google DeepMind; Google Research; Google Cloud; SNSF; IBM and Felicis.

### v2

@@ Abstract
The long-standing vision of general-purpose robots hinges on their ability to understand and act upon natural language instructions. Vision-Language-Action (VLA) models have made remarkable progress toward this goal, yet their generated actions can still misalign with the given instructions. In this paper, we investigate test-time verification as a means to shrink the “intention-action gap.” We first characterize the test-time scaling laws for embodied instruction following and demonstrate that jointly scaling the number of rephrased instructions and generated actions greatly increases test-time sample diversity, often recovering correct actions more efficiently than scaling each dimension independently. To capitalize on these scaling laws, we present CoVer, a contrastive verifier for vision–language–action alignment, and show that our architecture scales gracefully with additional computational resources and data. We then introduce CoVer-VLA, a hierarchical test-time verification pipeline using the trained verifier. At deployment, our framework precomputes a diverse set of rephrased instructions from a Vision-Language-Model (VLM), repeatedly generates action candidates for each instruction, and then uses the verifier to select the optimal high-level prompt and low-level action chunks. Compared to scaling policy pre-training on the same data, our verification approach yields 22% gains in-distribution and 13% out-of-distribution on the SIMPLER benchmark, with a further 45% improvement in real-world experiments. On the PolaRiS benchmark, CoVer-VLA achieves 14% gains in task progress and 9% in success rate.
@@ 1 Introduction
Figure 2: Test-Time Scaling Law for Embodied Instruction Following. Compared to prior methods that construct an action proposal distribution through repeated sampling [15] or Gaussian perturbations [12], we find that instruction rephrasing produces a broader set of action candidates, leading to improved recovery of the correct action. Furthermore, a hybrid test-time scaling strategy that increases both the number of rephrases and the number of sampled actions per rephrase is more effective than either strategy alone. We characterize each sampling approach using a power law, where the logarithm of oracle action error $e$ is a function of the number of action candidates $k$: $\log(e)\approx\log(a)+b\cdot\log(k)$.
@@ 4.1 Hierarchical Prompt-Action Optimization
Figure 3: Overview of CoVer Training Strategy. CoVer learns a joint embedding space aligning visual observations, language instructions, and robot actions through contrastive pre-training. A fused image–text encoder selectively extracts task-relevant visual features, while an action encoder projects action sequences into the same embedding space. This architecture enables cross-modal alignment between high-level instructions and executed behaviors.
@@ 5.4 Simulation Evaluation Results
Table 1: PolaRiS Evaluation Results. Mean task progress and success rate ($\pm$ standard deviation) across 50 episodes and 3 seeds on three PolaRiS environments using $\pi_{0.5}$ as the base robot policy. Pairing $\pi_{0.5}$ with CoVer consistently improves performance across all tasks, achieving a 13.9% gain in task progress and a 9.3% increase in success rate on average.
@@ 5.6 Latency Analysis and Optimizations
Figure 8: Real-World Evaluation Results. $\pi_{0}$ + CoVer significantly outperforms the baseline $\pi_{0}$ (rephrase), achieving a 45% absolute improvement in task success rate over the baseline policy.
@@ 6 Conclusion
In this paper, we present CoVer-VLA, a novel contrastive-based verifier and hierarchical test-time scaling framework that bridges the “intention–action gap” for generalist robot policies. CoVer-VLA achieves substantial performance improvements across both simulated and real-world settings, particularly under out-of-distribution conditions. Our findings demonstrate that allocating compute to reasoning and verification at deployment can be more effective than scaling policy training alone, providing a promising direction for robust policy deployment in the real world. While our study focuses on applying the verifier for test-time scaling, the same design and principle can extend beyond inference optimizations such as post-training with reinforcement learning or run-time monitoring. Future work could also explore more efficient architectures for both the base policy and verifier to further reduce latency and enable broader use of test-time scaling in real-world robotic settings.
@@ 7 Acknowledgments
We thank the members of the Stanford Autonomous Systems Lab, Scaling Intelligence Lab, and IRIS Lab for their constructive feedback and informative discussions. We gratefully acknowledge the support of DARPA; NASA ULI; Schmidt Sciences; Google DeepMind; Google Research; Google Cloud; SNSF; IBM and Felicis.